Could someone explain how to solve the following problem?
Please.
Find the accumulated amount after 5 years if $12,000 is invested at 18% per year compounded monthly.
Thank you!!!!

Answer by Alan3354(69443)

Find the accumulated amount after 5 years if $12,000 is invested at 18% per year compounded monthly.
-----------------
5 yrs = 60 months = # of periods
Interest rate/month = 0.18/12 = 0.015
FV = 12000*(1+0.015)^60
FV = 12000*2.44322
FV = $29318.64